OpenStudy (anonymous):

Give the tenth term in the geometric sequence. 2, 6, 18, 54, ... A.39,366 B. 13,122 C, 4,374 #2.What is the sixth term in this geometric sequence? 480, 240, 120, 60, ... A.15 B.30 C.7.5

Parth (parthkohli):

1. The formula is: \[\LARGE 2 * 3^{n - 1}\] Replace n with 10 and you get the answer. 2. The formula is: \[\LARGE 480/2^{n -1}\] Replace n with 6 and you get the answer.

OpenStudy (anonymous):

\[\LARGE a_n=a_1\cdot r^{n-1}\] and \(\LARGE r=\frac{a_n}{a_{n-1 }}\)